OpenStudy (anonymous):

use implicit differentiation to find dy/dx in terms of x and y 4x^2+4x^3y^2 = 3xy

hartnn (hartnn):

which step are you stuck ?

hartnn (hartnn):

x^3y^2 and xy will need product rule and chain rule. can u differentiate those ?

OpenStudy (anonymous):

confused on all of it mostly. i'm terrible at implicit differentiation

hartnn (hartnn):

i'll do x^3y^2 for you you try for xy (x^3y^2)' = x^3(y^2)' + y^2 (x^3)' = x^3 (2y dy/dx) + y^2 (3x^2) did u get this ?

hartnn (hartnn):

when you differentiate a function of y w.rt x , f(y) you get f'(y) dy/dx